On​ Monday, Harold picked up 4 donuts and 5 large coffees for the office staff. He paid 6.52  On​ Tuesday, Melinda picked up 5 donuts and 4 large coffees for the office staff. She paid 6.17  What is the cost of one​ donut? What is the cost of one large​ coffee?
Answer by Alan3354(69443)

On​ Monday, Harold picked up 4 donuts and 5 large coffees for the office staff. He paid 6.52  On​ Tuesday, Melinda picked up 5 donuts and 4 large coffees for the office staff. She paid 6.17  What is the cost of one​ donut? What is the cost of one large​ coffee?
======================================
4D + 5C = 652
5D + 4C = 617
-------------------------
Can you do the rest?
=========================
1 way to do this is using elimination.
Multiply eqn 1 by 5 and eqn 2 by 4.
--------------------
4D + 5C = 652 times 5 ---> 20D + 25C = 3260
5D + 4C = 617 times 4 ---> 20D + 16C = 2468
==========================
20D + 25C = 3260
20D + 16C = 2468
--------------------------- Subtract
9C = 792
C = 88
===========================
Sub for C in any eqn to find D.